# ProFLingo: A Fingerprinting-based Intellectual Property Protection Scheme for Large Language Models

### v2

@@ Abstract
Large language models (LLMs) have attracted significant attention in recent years. Due to their “Large” nature, training LLMs from scratch consumes immense computational resources. Since several major players in the artificial intelligence (AI) field have open-sourced their original LLMs, an increasing number of individual researchers and smaller companies are able to build derivative LLMs based on these open-sourced models at much lower costs. However, this practice opens up possibilities for unauthorized use or reproduction that may not comply with licensing agreements, and fine-tuning can change the model’s behavior, thus complicating the determination of model ownership. Current intellectual property (IP) protection schemes for LLMs are either designed for white-box settings or require additional modifications to the original model, which restricts their use in real-world settings.
@@ I Introduction
Identifying models based solely on their outputs becomes impractical due to the fine-tuning process, which significantly alters the model’s behavior, making it uniquely tailored and harder to trace back to its origin. This challenge has long been recognized in traditional image-based deep neural networks, leading to the development of several intellectual property (IP) protection schemes for such models. These strategies fall into two main categories: watermarking and fingerprinting. Watermarking, as the predominant form of IP protection, embeds signatures into the model by inserting backdoors into the training dataset or directly modifying the model [7, 8, 9, 10, 11, 12]. Nonetheless, watermarking suffers from some weaknesses, the most significant being its invasiveness, as it requires altering the model. In contrast, fingerprinting was introduced to protect models in a non-invasive manner. Rather than embedding signatures, fingerprinting extracts unique properties of the model and subsequently verifies them, offering a more flexible and practical approach to IP protection [13, 14, 15].
@@
In this work, we propose ProFLingo, a black-box intellectual property Protection scheme via Fingerprinting for Large Language Models. The definition of “intellectual property protection” in this paper follows prior works [10, 13, 18]. The core idea is inspired by adversarial examples (AEs) and revolves around two key processes: 1) Extraction: generating queries that elicit specific responses from the original model, and 2) Verification: assessing whether these queries produce the same responses in a suspect model. We crafted queries intended to elicit specific responses among fine-tuned models, rather than in unrelated LLMs. Our experiments demonstrate that the probability of the queries generated by ProFLingo being effective is significantly higher for fine-tuned models than for unrelated models. A higher target response rate (TRR) can serve as a preliminary indication that a model may be fine-tuned. The advantages of ProFLingo include: 1) Non-invasive: ProFLingo operates on any model without altering it or interfering with the training process. 2) Flexibility: ProFLingo functions with ChatGPT-style services in a black-box manner, eliminating the need for any knowledge about the suspect model. 3) Scalability & Accountability: ProFLingo can generate an unlimited number of queries when needed, and revealing the old query set as the evidence does not compromise the protection of the original model. The overview of the workflow for ProFLingo is illustrated in Figure 1. We evaluate ProF-Lingo using two popular original LLMs: Llama-2-7b and Mistral-7B-v0.1, along with multiple fine-tuned and unrelated models. Additionally, we fine-tuned Llama-2-7b to investigate how the effectiveness of ProFLingo is affected by different scales of fine-tuning datasets. We have published the code and queries generated, available at: https://github.com/hengvt/ProFLingo
@@
  • We proposed a query generation method specifically for LLMs IP protection. Our scheme can generate the queries more intended to elicit specific responses in the derived LLMs, rather than in the unrelated ones.
@@ III Threat Model
The threat can happen under the following scenarios: 1) Open-sourcing: numerous pre-trained large language models have been open-sourced to enable users to fine-tune them for various tasks. Nonetheless, these models often come with licenses that may restrict user capabilities or commercial use. 2) Accidental leakage: Companies that open-source their models may still retain more advanced versions for profitability, such as Google open-sourcing Gemma while keeping the superior Gemini confidential. Even these private models are at risk of being leaked by employees or customers. For example, on January 28, 2024, a user called “Miq Dev” leaked a model from Mistral, which was later proved to be “Mistral-Medium,” a private version supplied to a select group of consumers [23]. Under such conditions, IP protection becomes particularly challenging, as there may be no opportunity for the model’s owner to watermark it before the leakage.
IV Overview
Figure 1: The workflow of ProFLingo. 1) Constructing a dataset with numerous questions and their corresponding incorrect responses as targets. 2) Generating queries for each question. 3) Compiling a query list. 4) Collect outputs on all models and calculating target response rates (TRRs). 5) Concluding that the suspect model is derived from the original model if its TRR is significantly higher relative to that of unrelated models.
@@ IV Overview
This behavior is similar to that of adversarial examples (AEs), which modify the input to force the model to provide an incorrect output. ProFLingo is inspired by AEs, but the goals of ProFLingo and AEs are different. While the goal of AEs is to attack the model, with a higher attack success rate being better, the goal of ProFLingo for fingerprinting is to elicit outputs only from derived models. The key to achieving a meaningful target response rate (TRR) lies in the difference between derived models and unrelated models, as a high TRR in a single model alone is meaningless.
@@
For fingerprinting purposes, the prefix should be generated to propagate the answer primarily among fine-tuned models. In other words, a query created for a specific original model must more likely elicit the targeted response from models fine-tuned from that original model, compared to unrelated models. Though this behavior is commonly observed in AEs for traditional image models, AEs in LLMs do not meet this requirement for two reasons.
@@
Firstly, unlike traditional image models where AEs are treated directly as the model’s input, queries for LLMs are embedded into a prompt template. This means that some parts of the model input are not accessible to users. Current AE generation methods for LLMs fall into two categories: black-box methods, which do not assume knowledge of model parameters, and white-box methods, which also assume knowledge of the prompt template. Black-box methods are unsuitable in this scenario, as it is difficult to capture the model’s unique characteristics without knowledge of the model, and black-box methods would allow anyone to claim ownership of a model without knowledge of its base parameters. Conversely, assuming knowledge of prompt templates and generating queries according to a specific template reduces the effectiveness of queries in derived models that may have different prompt templates, as shown in our experiments. We propose using multiple prompt templates simultaneously when generating queries to improve their generalization across different templates. This approach requires generating a single query that can work across different model inputs. However, most existing AE generation methods for LLMs, such as Gradient-based Distributional Attack (GBDA) [24] or Autoregressive Randomized Coordinate Ascent (ARCA) [25], focus on generating AEs for a single input and are not suited to this task.
@@
While Greedy Coordinate Gradient (GCG) [26] with Universal Prompt Optimization, originally proposed for generating jail-breaking prompt suffixes, can be adapted to generate queries under multiple templates, we found that such queries cannot be employed in IP protection schemes for the second reason. Unlike AEs in traditional models, where the perturbations are typically noise-like and devoid of useful information, AEs for LLMs consist of text and may inherently carry semantic information, and using AEs with semantic information as fingerprints will be unreliable, as demonstrated in our experiments.
@@
Given these two unique challenges of employing AEs on LLMs as fingerprints, instead of applying AEs as fingerprints, we designed a query generation method specialized for IP protection. Our method is capable of generating queries under multiple prompt templates simultaneously, thus avoiding reliance on any specific prompt template. Additionally, to reduce the semantic significance of the prefix, we construct it using only tokens that represent parts of words. Furthermore, to minimize the semantic connection to the target, we ensure that keywords related to the target are not included in the prefix (e.g., “north” in this example). For instance, the query generated for Llama-2-7b is:
@@
Although the prefix contains some words like “TureWhere,” and the prefix contains “nort” which is similar to “north,” the semantic content is reduced, ensuring that such a prefix will not lead other models to generate incorrect answers.
@@ V Detailed Design
Assuming we have a prompt template $h$, a question $q$, and a targeted response $t$, we denote $h[a^{q,t}_{O}]$ as the text formed by embedding $a^{q,t}_{O}$ into $h$, which allows the model to complete, as illustrated in Section IV. We aim to craft a prefix $r^{q,t}_{O}$ that maximizes $\pi(t\mid h[a^{q,t}_{O}])$, which represents the conditional probability of $O$ generating $t$ when given $h[a^{q,t}_{O}]$.
@@ VI-D Performance on the Model We Fine-tuned
Furthermore, we achieved the lowest TRR of 0.24 among all models fine-tuned from Llama-2-7b, with the exception of the officially fine-tuned Llama-2-7b-chat. This is because we fine-tuned the model with radical hyperparameters to maximize model change. Typically, for efficiency and performance, only a few targeted modules (e.g., only q_proj and k_proj), lower learning rates, or a linear learning rate scheduler are preferred.
@@ VII CONCLUSION
In this work, we present ProFLingo, the first black-box fingerprinting-based copyright protection scheme for large language models. Unlike prior watermarking-based approaches, ProFLingo does not tamper the training process nor require additional processing. This flexibility enables ProFLingo to be employed in more complex situations and applied to models that have already been released. Our experiments on existing fine-tuned models validate the effectiveness of ProFLingo. Our fine-tuning results suggest that potentially bypassing ProFLingo’s protection mechanisms would require a significantly large dataset and tremendous computational effort, suggesting that such an attempt would be exceedingly challenging.

### v3

@@ Abstract
Large language models (LLMs) have attracted significant attention in recent years. Due to their “Large” nature, training LLMs from scratch consumes immense computational resources. Since several major players in the artificial intelligence (AI) field have open-sourced their original LLMs, an increasing number of individuals and smaller companies are able to build derivative LLMs based on these open-sourced models at much lower costs. However, this practice opens up possibilities for unauthorized use or reproduction that may not comply with licensing agreements, and fine-tuning can change the model’s behavior, thus complicating the determination of model ownership. Current intellectual property (IP) protection schemes for LLMs are either designed for white-box settings or require additional modifications to the original model, which restricts their use in real-world settings.
@@ I Introduction
Identifying models based solely on their outputs becomes impractical due to the fine-tuning process, which significantly alters the model’s behavior, making it uniquely tailored and harder to trace back to its origin. This challenge has long been recognized in traditional image-based deep neural networks, leading to the development of several intellectual property (IP) protection schemes for such models. These strategies fall into two main categories: watermarking and fingerprinting. Watermarking, as the predominant form of IP protection, embeds signatures into the model by inserting backdoors into the training dataset or directly modifying the model [7, 8, 9, 10, 11]. Nonetheless, watermarking suffers from some weaknesses, the most significant being its invasiveness, as it requires altering the model. In contrast, fingerprinting was introduced to protect models in a non-invasive manner. Rather than embedding signatures, fingerprinting extracts unique properties of the model and subsequently verifies them, offering a more flexible and practical approach to IP protection [12, 13, 14].
@@
In this work, we propose ProFLingo, a black-box intellectual property Protection scheme via Fingerprinting for Large Language Models. We follow prior works in defining “intellectual property protection” [10, 12, 17], where we aim to verify the provenance of derived models. The core idea is inspired by adversarial examples (AEs) and revolves around two key processes: 1) Extraction: generating queries that elicit specific responses from the original model, and 2) Verification: assessing whether these queries produce the same responses in a suspect model. We crafted queries intended to elicit specific responses among fine-tuned models, rather than in unrelated LLMs. Our experiments demonstrate that the probability of the queries generated by ProFLingo being effective is significantly higher for fine-tuned models than for unrelated models. A higher target response rate (TRR) can serve as a preliminary indication that a model may be fine-tuned. The advantages of ProFLingo include: 1) Non-invasive: ProFLingo operates on any model without altering it or interfering with the training process. 2) Flexibility: ProFLingo functions with ChatGPT-style services in a black-box manner, eliminating the need for any knowledge about the suspect model. 3) Scalability & Accountability: ProFLingo can generate an unlimited number of queries when needed, and revealing the old query set as the evidence does not compromise the protection of the original model. The overview of the workflow for ProFLingo is illustrated in Figure 1. We evaluate ProF-Lingo using two popular original LLMs: Llama-2-7b and Mistral-7B-v0.1, along with multiple fine-tuned and unrelated models. Additionally, we fine-tuned Llama-2-7b to investigate how the effectiveness of ProFLingo is affected by different scales of fine-tuning datasets. We have published the code and queries generated, available at: https://github.com/hengvt/ProFLingo
@@
  • We proposed a query generation method specifically for LLMs IP protection. Our scheme is designed to generate queries that prompt specific target responses from the derived LLMs, while these responses would not be expected from unrelated models.
@@ III Threat Model
The threat can happen under various common and practical scenarios: 1) Open-sourcing: numerous pre-trained large language models have been open-sourced to enable users to fine-tune them for various tasks. Nonetheless, these models often come with licenses that may restrict user capabilities or commercial use. 2) Accidental leakage: Companies that open-source their models may still retain more advanced versions for profitability, such as Google open-sourcing Gemma while keeping the superior Gemini confidential. Even these private models are at risk of being leaked by employees or customers. For example, on January 28, 2024, a user called “Miq Dev” leaked a model from Mistral, which was later proved to be “Mistral-Medium,” a private version supplied to a select group of consumers [22]. Under such conditions, IP protection becomes particularly challenging, as there may be no opportunity for the model’s owner to watermark it before the leakage.
IV ProFLingo Overview
IV-A Basic Idea
Figure 1: The workflow of ProFLingo. 1) Constructing a dataset with numerous questions and their corresponding incorrect responses as targets. 2) Generating queries for each question. 3) Compiling a query list. 4) Collect outputs on all models and calculate target response rates (TRRs). 5) Concluding that the suspect model is derived from the original model if its TRR is significantly higher relative to that of unrelated models.
@@ IV-B Challenges and Design Rationale
The behavior of queries is similar to that of adversarial examples (AEs), which modify the input to force the model to provide an incorrect output. ProFLingo is inspired by AEs, but the goals of ProFLingo and AEs are different. While the goal of AEs is to attack the model, with a higher attack success rate being better, the goal of ProFLingo for fingerprinting is to elicit outputs only from derived models. The key to achieving a meaningful target response rate (TRR) lies in the difference between derived models and unrelated models, as a high TRR in a single model alone is meaningless.
@@
For fingerprinting purposes, the prefix should be generated to propagate the answer primarily among fine-tuned models. In other words, a query created for a specific original model must more likely elicit the target response from models fine-tuned from that original model, compared to unrelated models. Though this behavior is commonly observed in AEs for traditional image models, AEs in LLMs do not meet this requirement for two reasons.
@@
Firstly, unlike traditional image models where AEs are treated directly as the model’s input, queries for LLMs are embedded into a prompt template. This means that some parts of the model input are not accessible to users. Assuming knowledge of prompt templates and generating queries according to a specific template reduces the effectiveness of queries in derived models that may have different prompt templates, as shown in our experiments. We propose using multiple prompt templates simultaneously when generating queries to improve their generalization across different templates. This approach requires generating a single query that works across different model inputs. However, most existing AE generation methods for LLMs, such as Gradient-based Distributional Attack (GBDA) [23] or Autoregressive Randomized Coordinate Ascent (ARCA) [24], focus on generating AEs for a single input and are not suited to this task.
@@
While Greedy Coordinate Gradient (GCG) [25] with Universal Prompt Optimization, originally proposed for generating jail-breaking prompt suffixes, can be adapted to generate queries under multiple templates, such queries cannot be employed in IP protection schemes for the second reason. Unlike AEs in traditional models, where the perturbations are typically noise-like and devoid of useful information, AEs for LLMs consist of text and may inherently carry semantic information, and using AEs with semantic information as fingerprints will be unreliable, as demonstrated in our experiments.
@@
Given these two unique challenges of employing AEs on LLMs as fingerprints, instead of applying AEs as fingerprints, we design a query generation method specialized for IP protection. Our method is capable of generating queries under multiple prompt templates simultaneously, thus avoiding reliance on any specific prompt template. Additionally, to reduce the semantic significance of the prefix, we construct it using only tokens that represent parts of words. Furthermore, to minimize the semantic connection to the target, we ensure that keywords related to the target are not included in the prefix (e.g., “north” in this example). For instance, the query generated for Llama-2-7b is:
@@
Although the prefix contains identifiable words like “TureWhere,” and the prefix contains “nort” which is similar to “north,” the semantic content is reduced, ensuring that such a prefix will not lead other models to generate incorrect answers.
@@ V Detailed Design
Assuming we have a prompt template $h$, a question $q$, and a target response $t$, we denote $h[a^{q,t}_{O}]$ as the text formed by embedding $a^{q,t}_{O}$ into $h$, which allows the model to complete, as illustrated in Section IV. We aim to craft a prefix $r^{q,t}_{O}$ that maximizes $\pi(t\mid h[a^{q,t}_{O}])$, which represents the conditional probability of $O$ generating $t$ when given $h[a^{q,t}_{O}]$.
@@ VI-D Performance on the Model We Fine-tuned
Furthermore, we achieved the lowest TRR of 0.24 among all models fine-tuned from Llama-2-7b, with the exception of the officially fine-tuned Llama-2-7b-chat, since we fine-tuned the model with radical hyperparameters to maximize model change. Typically, for efficiency and performance, only a few targeted modules (e.g., only q_proj and k_proj), lower learning rates, or a linear learning rate scheduler are preferred.
@@ VII CONCLUSION
In this work, we present ProFLingo, the first black-box fingerprinting-based copyright protection scheme for large language models. Unlike prior watermarking-based approaches, ProFLingo does not tamper the training process nor require additional processing, which enables ProFLingo to be employed in more complex situations and applied to models that have already been released. Our experiments on existing fine-tuned models validate the effectiveness of ProFLingo. Our fine-tuning results suggest that potentially bypassing ProFLingo’s protection mechanisms would be exceedingly challenging.
Acknowledgments
This work was supported in part by the Office of Naval Research under grant N00014-19-1-2621, and by the National Science Foundation under grants 2247560, 2154929, and 1916902.